Question 10 of 22 A 4.05 g sample of an unknown gas at 35 °C and 1.05 atm is stored in a 2.15 L flask. What is the density ofQuestion 14 of 22 A gas mixture is made by combining 6.7 g each of Ar, Ne, and an unknown diatomic gas. At STP, the mixture o< Question 6 of 22 > If 48.5 mol of an ideal gas occupies 97.5 L at 51.00 °C, what is the pressure of the gas? P= atm< Question 1 of 22 > Perform the conversions. 547.9 mmHg = atm 1.087 atm = Torr 328.0 kPa = atm 647.1 mmHg = Pa
Answer #1

PV = nRT 145 gun/mole molan mall W-4.05gm Atomic mass of f = 18.998 T=35°C T=35+273=308K : fe has molan mass of P=1.05 atm =